# Knowledge Transfer from High-Resource to Low-Resource Programming Languages for Code LLMs

### v5

@@ Abstract.
Over the past few years, Large Language Models of Code (Code LLMs) have started to have a significant impact on programming practice. Code LLMs are also emerging as building blocks for research in programming languages and software engineering. However, the quality of code produced by a Code LLM varies significantly by programming language. Code LLMs produce impressive results on programming languages that are well represented in their training data (e.g., Java, Python, or JavaScript), but struggle with low-resource languages that have limited training data available. Low resource languages include OCaml, Racket, and several others.
This paper presents an effective approach for boosting the performance of Code LLMs on low-resource languages using semi-synthetic data. Our approach generates high-quality datasets for low-resource languages, which can then be used to fine-tune any pretrained Code LLM. Our approach, called MultiPL-T, translates training data from high-resource languages into training data for low-resource languages in the following way. 1) We use a Code LLM to synthesize tests for commented code from a high-resource language, filtering out faulty tests and code with low test coverage. 2) We use a Code LLM to translate Python code to a target low-resource language, and use tests to validate the translation. We apply this approach to generate tens of thousands of new, validated training items for Julia, Lua, OCaml, R, and Racket. Furthermore, we use an open model (StarCoderBase) with open training data (The Stack), which allows us to decontaminate benchmarks, train models without violating licenses, and run experiments that could not otherwise be done.
With MultiPL-T generated data, we present fine-tuned versions of StarCoderBase and Code Llama for Julia, Lua, OCaml, R, and Racket. On established benchmarks (MultiPL-E), these models outperform other open Code LLMs. The MultiPL-T approach is easy to apply to new languages, and is significantly more efficient and effective than alternatives such as training longer.
copyright: none
@@ 1. Introduction
Large Language Models of Code (Code LLMs) are starting to have a significant impact on both professional programmers and research in programming languages and software engineering. GitHub Copilot is just one of several popular tools powered by Code LLMs (Copilot, 2023; CodeWhisperer, 2023; TabNine, 2023). Moreover, Code LLMs are also emerging as a building block for research (Schäfer et al., 2023; Bareiß et al., 2022; Murali et al., 2023; Nam et al., 2023; Ross et al., 2023; Lemieux et al., 2023; Xia et al., 2023; Joshi et al., 2023; Phung et al., 2023; First et al., 2023; Chen et al., 2023). However, the quality of code produced by a Code LLM varies significantly by programming language. Models are most impressive at high-resource programming languages such as Python, JavaScript, and Java, but struggle to produce code in low-resource languages, such as Racket and OCaml (Cassano et al., 2023; Athiwaratkun et al., 2022; Zheng et al., 2023). This puts programmers who rely on these languages at a disadvantage, since they do not receive the same benefits that Code LLMs can deliver for high-resource languages (Ziegler et al., 2022; Murali et al., 2023).
@@
The key issue is that the performance of Code LLMs depends on the amount of language data available for training. For example, The Stack, which is the training set for several contemporary Code LLMs, has 64GB of Python, but only around 1GB of OCaml and 0.5GB of Scheme/Racket (Kocetkov et al., 2023). [^1] This is the volume of data that remains after files are deduplication for training (Li et al., 2023). As figure 1 shows, the performance of StarCoderBase, an open Code LLM trained on The Stack, generally increases as the proportion of training data for the language increases.
@@
Figure 1. The performance of StarCoderBase-15B on several languages from the MultiPL-E benchmark for Code LLMs, plotted against their proportion of the model’s training data. Using MultiPL-T, this paper significantly improves how StarCoderBase-15B performs on several low-resource languages, as shown by the arrows. The bottom of each arrow indicates how the base model performs, and the arrowheads indicate performance after fine-tuning with MultiPL-T.
@@
Our goal in this paper is to investigate methods to further train, or fine-tune, pretrained Code LLMs to improve their performance on low-resource languages. The obvious approach is try to find more data, but for a low-resource language, more data is hard to find by definition. For example, The Stack already includes all permissively licensed code for 358 programming languages from GitHub as of 2022, and GitHub is by far the largest repository of open source code (Borges et al., 2016). [^2] The Stack deliberately excludes copyleft licenses and unlicensed code. An alternative is to train longer (i.e, for more epochs) on existing data. However, in this paper we show that training longer on several low-resource languages is not only inefficient, but can actually hurt performance (§ 3.1). Another alternative is to train models on synthetic data. Self-instruction (Wang et al., 2023) and instruction evolution (Luo et al., 2023) fine-tune LLMs on data generated by another LLM. These approaches have been effective at fine-tuning models on high-resource programming languages. However, we show that self-instruction does not work for low-resource languages, because for low-resource languages, the LLM-generated programs are of poor quality (§ 3.2). Here, ‘quality’ refers to both the correctness of the code (i.e., its ability to function as intended) and its stylistic integrity (i.e., adherence to best practices and readability).
@@
In this paper, we present a new and effective approach for fine-tuning Code LLMs for low-resource programming languages that is based on generating semi-synthetic training data. Our approach relies on several key ingredients. 1) The large volume of training data for high-resource programming languages includes a lot of well-documented code; 2) Code LLMs are effective and efficient unit test generators, and we can check that generated tests are valid (Schäfer et al., 2023); 3) We can mechanically translate many unit tests to a low-resource language with a simple compiler (Roziere et al., 2021; Athiwaratkun et al., 2022; Cassano et al., 2023); 4) Code LLMs can translate code from one language to another, and we can test these translations with the aforementioned tests, retry until tests pass, and engineer a prompt to increase the likelihood of a successful translation. Putting these four ideas together, we develop a pipeline for transferring training data across multiple programming languages that we call MultiPL-T.
@@
Using training data generated by MultiPL-T, we present fine-tuned Code LLMs that achieve state-of-the-art performance on five low-resource languages: Racket, OCaml, Lua, R, and Julia. We focus primarily on fine-tuning the StarCoder family of Code LLMs (Li et al., 2023). There are StarCoder models available at a variety of sizes, including a 1B parameter model that is lightweight enough to run on CPUs, and a more capable 15B parameter model, which we use as the test generator and language translator for MultiPL-T. The StarCoder models also have open training data, which allows us to compare MultiPL-T to a baseline of training longer on existing data for low-resource languages. Finally, we also present fine-tuned versions of the Code Llama 34B (Rozière et al., 2023) model, which was recently released.
@@
    Efficient fine-tuning datasets for Julia, Lua, OCaml, R, and Racket, comprising tens of thousands of documented and tested functions.
@@
    Fine-tuned versions of StarCoderBase and Code Llama for Julia, Lua, OCaml, R, and Racket in three sizes: 1B, 15B, and 34B parameters. The 1B models can be run relatively well on CPUs, whereas the larger models require GPUs to produce results with reasonable latency.
@@
    A thorough evaluation that includes a) a comparison of MultiPL-T to the baseline of training further on existing data, b) an evaluation of the fine-tuning efficiency with MultiPL-T, c) results on prior multi-language benchmarks (Cassano et al., 2023), d) a new multi-language benchmark designed to exercise in-context learning, e) an evaluation of how generated code adheres to common Racket programming style.
The MultiPL-T datasets, models, and the code to reproduce our work are available at:
https://huggingface.co/datasets/nuprl/MultiPL-T
@@ 2. Background
In this section, we give a high-level overview of how Code LLMs are trained and evaluated. We use StarCoder as the example, since it is the model that we use for most of MultiPL-T.
Training Large Language Models of Code
@@
The only way to build a training set of this scale is to scrape public repositories of code. There are a handful of public training sets that are based on GitHub (Xu et al., 2022; Felipe Hoffa, 2016), but the largest and most recent is The Stack (Kocetkov et al., 2023). The Stack v1.2 has 3TB of permissively licensed source code for 358 programming languages. It was constructed in 2022, and has since been used to train several Code LLMs (Allal et al., 2023; Replit, 2023; Nijkamp et al., 2023), including StarCoderBase. Specifically, StarCoderBase was trained on a subset of The Stack consisting of 86 programming languages.
@@
    a’, e’, i’, o’, u’. Here, y is also a vowel, but only when it is at the
@@
   a’, e’, i’, o’, u’. Here, y is also a vowel, but only when it is at the
@@
StarCoder is a family of models that are available at several sizes (Li et al., 2023). The largest and most capable model is the family is called StarCoderBase: it has 15B parameters, which requires a capable GPU to run inference with reasonable latency. There are smaller versions of StarCoderBase that were trained on exactly the same data. To make use of limited GPU resources, we use the smallest model, StarCoderBase-1B, for most experiments in this paper. However, we also show that our results generalize to StarCoderBase-15B. There is also a model in the StarCoder family that is just named StarCoder: it is StarCoderBase-15B specialized to excel at Python. [^3] StarCoder fine-tunes StarCoderBase-15B on two more epochs of Python data from The Stack. This paper uses StarCoderBase-15B for translations to low-resource languages, and StarCoder-15B for Python test generation.
@@
The Code Llama (Rozière et al., 2023) family of models were recently released and perform better than the StarCoder models on common benchmarks. While the authors state that the training data comes from publicly accessible datasets, they do not disclose the specific datasets used, preventing us from conducting the exhaustive evaluation on Code Llama that we do with StarCoder and its training data. Moreover, the Llama license forbids using model outputs to train non-Llama models, which is why MultiPL-T uses StarCoder for data generation. However, we train and evaluate the largest Code Llama on data constructed with StarCoder models.
@@
After training, a model can be further trained, or fine-tuned, with significantly fewer resources. For example, there are several fine-tuned versions of StarCoderBase that were trained with a few days of GPU time on a modest amount of data (e.g., (Luo et al., 2023; Muennighoff et al., 2023)). Most fine-tuned versions of StarCoderBase are designed to make the model even better at high-resource languages, such as Python. In contrast, this paper presents fine-tuned versions of StarCoderBase that are significantly better at several low-resource languages.
Benchmarking Code LLMs
Most Code LLM benchmarks, including those we use in this paper, follow the format introduced by the Codex “HumanEval” benchmark (Chen et al., 2021). Every benchmark problem has two parts: 1) a prompt for the LLM that has a function signature and a comment, and 2) a suite of test cases that are not given to the LLM. Thus each problem is run in two steps: 1) the model generates a function from the prompt, and 2) the generated function is then tested with the hidden tests, and all tests must pass for the generated code to be considered correct.
@@
Code LLMs appear to produce higher quality code when their output is sampled (Chen et al., 2021). Since sampling introduces non-determinism, we must evaluate their output by generating several samples from the same prompt. The most widely used metric for Code LLM performance is pass@k, which is the likelihood that the LLM produces a correct program at least once from $k$ attempts. Pass@k must be estimated from $n>>k$ samples. When $k=1$, pass@1 is the same as the pass rate ($k/n$). We use pass@1 as the metric for all our benchmarking experiments, which is common practice. Intuitively, pass@1 measures the ability of the Code LLM to generate a correct solution in a single attempt.

### v6

@@ Abstract.
Over the past few years, Large Language Models of Code (Code LLMs) have started to have a significant impact on programming practice. Code LLMs are also emerging as building blocks for research in programming languages and software engineering. However, the quality of code produced by a Code LLM varies significantly by programming language. Code LLMs produce impressive results on high-resource programming languages that are well represented in their training data (e.g., Java, Python, or JavaScript), but struggle with low-resource languages that have limited training data available (e.g., OCaml, Racket, and several others).
This paper presents an effective approach for boosting the performance of Code LLMs on low-resource languages using semi-synthetic data. Our approach, called MultiPL-T, generates high-quality datasets for low-resource languages, which can then be used to fine-tune any pretrained Code LLM. MultiPL-T translates training data from high-resource languages into training data for low-resource languages in the following way. 1) We use a Code LLM to synthesize unit tests for commented code from a high-resource source language, filtering out faulty tests and code with low test coverage. 2) We use a Code LLM to translate the code from the high-resource source language to a target low-resource language. This gives us a corpus of candidate training data in the target language, but many of these translations are wrong. 3) We use a lightweight compiler to compile the test cases generated in (1) from the source language to the target language, which allows us to filter our obviously wrong translations. The result is a training corpus in the target low-resource language where all items have been validated with test cases. We apply this approach to generate tens of thousands of new, validated training items for five low-resource languages: Julia, Lua, OCaml, R, and Racket, using Python as the source high-resource language. Furthermore, we use an open Code LLM (StarCoderBase) with open training data (The Stack), which allows us to decontaminate benchmarks, train models without violating licenses, and run experiments that could not otherwise be done.
Using datasets generated with MultiPL-T, we present fine-tuned versions of StarCoderBase and Code Llama for Julia, Lua, OCaml, R, and Racket that outperform other fine-tunes of these base models on the natural language to code task. We also present Racket fine-tunes for two very recent models, DeepSeek Coder and StarCoder2, to show that MultiPL-T continues to outperform other fine-tuning approaches for low-resource languages. The MultiPL-T approach is easy to apply to new languages, and is significantly more efficient and effective than alternatives such as training longer.
Large Language Models trained on Code
copyright: rightsretaineddoi: 10.1145/3689735journalyear: 2024journal: PACMPLjournalvolume: 8journalnumber: OOPSLA2article: 295publicationmonth: 10submissionid: oopslab24main-p302-pccs: Software and its engineering Automatic programming
@@ 1. Introduction
Large Language Models of Code (Code LLMs) are starting to have a significant impact on both professional programmers and research in programming languages and software engineering. GitHub Copilot is just one of several popular tools powered by Code LLMs (Copilot, 2023; CodeWhisperer, 2023; TabNine, 2023). Moreover, Code LLMs are also emerging as a building block for research (Schäfer et al., 2024; Bareiß et al., 2022; Murali et al., 2023; Nam et al., 2024; Ross et al., 2023; Lemieux et al., 2023; Xia et al., 2024; Joshi et al., 2023; Phung et al., 2023; First et al., 2023; Chen et al., 2023). However, the quality of code produced by a Code LLM varies significantly by programming language. Models are most impressive at producing code in high-resource programming languages such as Python, JavaScript, and Java, but struggle in low-resource languages, such as Racket and OCaml (Cassano et al., 2023; Athiwaratkun et al., 2022; Zheng et al., 2023). This puts programmers who rely on these languages at a disadvantage, since they do not receive the same benefits that Code LLMs can deliver for high-resource languages (Ziegler et al., 2022; Murali et al., 2023).
@@
The key issue is that the performance of Code LLMs depends on the amount of language data available for training. For example, The Stack, which is the training set for several contemporary Code LLMs, has 64GB of Python, but only around 1GB of OCaml and 0.5GB of Scheme/Racket (Kocetkov et al., 2023). [^1] This is the volume of data that remains after files are deduplication for training (Li et al., 2023). As figure 1 shows, the performance of StarCoderBase, an open Code LLM trained on The Stack, generally increases as the training data for the language increases.
@@
Figure 1. The performance of StarCoderBase-15B on several languages supported by the MultiPL-E benchmark for Code LLMs, plotted against their proportion of the model’s training data. Using MultiPL-T, this paper significantly improves how StarCoderBase-15B performs on several low-resource languages, as shown by the arrows. The bottom of each arrow indicates how the base model performs, and the arrowheads indicate performance after fine-tuning with MultiPL-T. We also show significant improvement on other LLMs (§ 5).
[…]
@@
Our goal in this paper is to investigate methods to further train, or fine-tune, pretrained Code LLMs to improve their performance on low-resource languages. The obvious approach is to try to find more data, but for a low-resource language, more data is hard to find by definition. For example, The Stack already includes all permissively licensed code for 358 programming languages from GitHub as of 2022, and GitHub is by far the largest repository of open-source code (Borges et al., 2016). [^2] The Stack deliberately excludes copyleft licenses and unlicensed code. An alternative is to train longer (i.e, for more epochs) on existing data. However, in this paper, we show that training longer on several low-resource languages is not only inefficient, but can actually hurt performance (§ 3.1). Another alternative is to train models on synthetic data, which is data generated by an LLM itself. These synthetic fine-tuning datasets are effective for high-resource programming languages (Wang et al., 2023; Luo et al., 2024). However, we show that synthetic data does not work for low-resource languages for the intuitive reason that LLMs generate poor quality programs in low-resource languages (§ 3.2).
Our approach
@@
In this paper, we present a new and effective approach for fine-tuning Code LLMs for low-resource programming languages that is based on generating semi-synthetic training data. Our approach relies on several key ingredients. 1) The large volume of training data for high-resource programming languages includes a lot of well-documented code; 2) Code LLMs are effective and efficient unit test generators, and we can check that generated tests are valid (Schäfer et al., 2024); 3) We can compile many unit tests to a low-resource language with a simple compiler (Roziere et al., 2021; Athiwaratkun et al., 2022; Cassano et al., 2023); 4) Code LLMs can translate code from one language to another, and although these translations may be faulty, we can filter them with the aforementioned tests, retry until tests pass, and engineer a prompt to increase the likelihood of a successful translation. Putting these four ideas together, we develop a pipeline for transferring training data across multiple programming languages that we call MultiPL-T.
[…]
See caption.
@@
Using training data generated by MultiPL-T, we present fine-tuned Code LLMs that achieve state-of-the-art performance on five low-resource languages: Racket, OCaml, Lua, R, and Julia. We focus primarily on fine-tuning the StarCoder family of Code LLMs (Li et al., 2023). There are StarCoder models available in a variety of sizes, including a 1B parameter model that is lightweight enough to run on CPUs, and a more capable 15B parameter model, which we use as the test generator and language translator for MultiPL-T. The StarCoder models also have open training data, which allows us to compare MultiPL-T to a baseline of training longer on existing data for low-resource languages. We also present fine-tuned versions of the Code Llama 34B and 70B (Rozière et al., 2023) models, and Racket fine-tunes of the recently released DeepSeek Coder (Guo et al., 2024) and StarCoder2 models (Lozhkov et al., 2024).
@@
    Efficient fine-tuning datasets for Julia, Lua, OCaml, R, and Racket, comprising tens of thousands of documented and tested functions generated with StarCoderBase-15B.
@@
    Fine-tuned versions of StarCoderBase 1B and 15B for Julia, Lua, OCaml, R, and Racket. For these languages, these fine-tuned models outperform prior fine-tunes of StarCoderBase on the natural language to code task.
@@
    Fine-tuned versions of Code Llama 34B and 70B for Julia, Lua, OCaml, R, and Racket. For these languages, these fine-tuned models outperform prior fine-tunes of Code Llama. More significantly, this is an uncommon result where data generated from a smaller model (StarCoderBase-15B) improves the performance of larger and better models (Code Llama 34B and 70B).
  6. (6)
    Fine-tuned versions of StarCoder2-15B and DeepSeek Coder 33B for Racket that also outperform other fine-tuned models. These are two very recently released models.
  7. (7)
    A thorough evaluation that includes a) a comparison of MultiPL-T to the baseline of training further on existing data, b) an evaluation of the fine-tuning efficiency with MultiPL-T, c) results on prior multi-language benchmarks (Cassano et al., 2023), d) a new multi-language benchmark designed to exercise in-context learning, e) an evaluation of how generated code adheres to the common Racket programming style, f) the impact of data deduplication, and g) the impact of fine-tuning on the Python source data.
@@ 2. Background
In this section, we give a high-level overview of how Code LLMs are trained and evaluated. We use StarCoder as the example, since it is the model that we use for most of our work.
2.1. Training and Fine-Tuning Large Language Models of Code
@@
The only way to build a training set of this scale is to scrape public repositories of code. There are a handful of public training sets that are based on GitHub (Xu et al., 2022; Felipe Hoffa, 2016), and The Stack (Kocetkov et al., 2023) is a recent example. The Stack v1.2 has 3TB of permissively licensed source code for 358 programming languages. It was constructed in 2022, and has since been used to train several Code LLMs (Allal et al., 2023; Replit, 2023; Nijkamp et al., 2023), including StarCoderBase. Specifically, StarCoderBase was trained on a filtered subset of The Stack consisting of 86 programming languages.
@@
’a’, ’e’, ’i’, ’o’, ’u’. Here, ’y’ is also a vowel, but only when it is at the
@@
’a’, ’e’, ’i’, ’o’, ’u’. Here, ’y’ is also a vowel, but only when it is at the
@@
StarCoder is a family of models that are available at several sizes (Li et al., 2023). The largest and most capable model in the family is called StarCoderBase, which has 15B parameters. There are smaller versions of StarCoderBase that were trained on exactly the same data. To make use of limited GPU resources, we use the smallest model, StarCoderBase-1B, for most experiments in this paper. However, we also show that our results generalize to StarCoderBase-15B. There is also a model in the StarCoder family that is just named StarCoder: it is StarCoderBase-15B specialized to excel at Python. [^3] StarCoder fine-tunes StarCoderBase-15B on two more epochs of Python data from The Stack. This paper uses StarCoderBase-15B for translations to low-resource languages, and StarCoder-15B for Python test generation.
@@
The Code Llama (Rozière et al., 2023) family of models were recently released and perform better than the StarCoder models on common benchmarks. While the authors state that the training data comes from publicly accessible datasets, they do not disclose the specific datasets used, preventing us from conducting the exhaustive evaluation on Code Llama that we do with StarCoder and its training data. Moreover, the Llama license forbids using model outputs to train non-Llama models, which is why we use StarCoder for data generation. However, we train and evaluate the larger Code Llama models (34B and 70B).
@@
After training, a model can be further trained, or fine-tuned, with significantly fewer resources. For example, there are several fine-tuned versions of StarCoderBase that were trained with a few days of GPU time on a modest amount of data (e.g., (Luo et al., 2024; Muennighoff et al., 2024)). Most fine-tuned versions of StarCoderBase are designed to make the model even better at high-resource languages, such as Python. In contrast, this paper presents fine-tuned versions of StarCoderBase that are significantly better at several low-resource languages.
It is common to distill data from a larger model (e.g., GPT-4), to fine-tune a smaller model (Luo et al., 2024; Wei et al., 2024; Gunasekar et al., 2023). However, we show that MultiPL-T can do the reverse: we use data generated from StarCoderBase-15B to fine-tune larger models, CodeLlama-34B and 70B. This is a form of weak-to-strong supervision (Burns et al., 2024), where a smaller model is used to generate data for training a larger model, showing the scalability of MultiPL-T.
2.2. Code LLM Tasks and Benchmarking Code LLMs
A Code LLM can be prompted to perform a wide variety of tasks, including code translation (e.g., (Pan et al., 2024)), test generation (e.g., (Schäfer et al., 2024)), code mutation (e.g., (Xia et al., 2024)), code editing (e.g., (Cassano et al., 2024)), and much more. This article focuses on the natural language to code task (e.g., (Heidorn, 1974)) for low-resource programming languages. Making Code LLMs better on other tasks is beyond the scope of this article.
Most Code LLM benchmarks for the natural language to code task, including those we use in this paper, follow the format introduced by the Codex “HumanEval” benchmark (Chen et al., 2021). Every benchmark problem has two parts: 1) a prompt for the LLM that has a function signature and a comment, and 2) a suite of test cases that are not given to the LLM. Thus each problem is run in two steps: 1) the LLM generates a function from the prompt, and 2) the generated function is then tested with the hidden tests, and all tests must pass for the generated code to be considered correct.
@@ 2.2. Code LLM Tasks and Benchmarking Code LLMs
Code LLMs appear to produce higher-quality code when their output is sampled (Chen et al., 2021). Since sampling introduces non-determinism, we must evaluate their output by generating several samples from the same prompt. The most widely used metric for Code LLM performance is pass@k, which is the likelihood that the LLM produces a program that passes all hidden tests at least once from $k$ attempts. Pass@k must be estimated from $n>>k$ samples. When $k=1$ and there are $c$ successes, pass@1 is the same as the pass rate ($c/n$). We use pass@1 as the metric for all our benchmarking experiments, which is common practice. Intuitively, pass@1 measures the ability of the Code LLM to generate a correct solution in a single attempt.
2.3. Why StarCoder?
In the rest of this paper, the majority of the work that we present uses the StarCoder family of Code LLMs for the following reasons.
  1. (1)
    At the time that we started this work, StarCoder was the best-performing, permissively licensed, open Code LLM available.
  2. (2)
    The StarCoder family includes a fairly small 1B parameter model, which is amenable to experiments on a budget.
  3. (3)
    Although there were better closed-sourced LLMs when we started, they (a) did not support fine-tuning, (b) were significantly more expensive to use at the scale of our work, or (c) prohibited using their outputs to fine-tune other LLMs (OpenAI, 2023b; Google, 2023; Anthropic, 2023b).
  4. (4)
    StarCoder remains the only Code LLM with open training data, which we use to correlate model performance with training set size (figure 1), evaluate further training (§ 3.1), and to check that our benchmarks are not in the training data. It would not have been possible to do this work with any other Code LLM.
We do fine-tune several newer Code LLMs with StarCoder-generated data to show that our approach remains useful (§ 5).